. A granite wall has a thickness of 0.50 m and a thermal conductivity of 2.1 W/(m · C°). The temperature on one face of the wall is 5.2 °C and 20.0 °C on the opposite face. How much heat is transferred in one hour through each square meter of the granite wall?
Answer #1

use conduction formula:

Q/t = KA*(T2-T1)/d
Q/3600 = 2.1*1*(20.0 - 5.2)/0.50
Q /3600 = 62.16
Q = 223776 J

Answer: 223776 J